Jane Cooper's Department Stores allocates the costs of the Personnel and Payroll departments to three retail sales departments, Housewares, Clothing, and Furniture. In addition to providing services to the operating departments, Personnel and Payroll provide services to each other. Copper's allocates Personnel Department costs on the basis of the number of employees and Payroll Department costs on the basis of gross payroll. Cost and allocation information for June is as follows:

Personnel Payroll Houseware Clothing Furniture
Direct Department Cost $15,600 $6,400 $24,400 $40,000 $33,500
Number of Employees 15 12 24 32 12
Gross Payroll $12,000 $6,600 $21,200 $26,100 $16,200

(a) Determine the percentage of total Personnel Department services that was provided to the Payroll Department.

(b) Determine the percentage of total Payroll Department services that was provided to the Personnel Department.

(c) Prepare a schedule showing Personnel Department and Payroll Department cost allocations to the operating departments, assuming Cooper's uses the step method.
Answer #1

@ Page, Alo 0 Perceologe provided to the geysel department Total employees = 15+12 + 24 + 32+12 7. provided to the payroll de1 Page No @ © Schedule of cof allocations Personnel department could be allocated to pay oroll, howeitores clothing a furnituPage No ③ Amount of gross paysroll of 3 department - 01200 +26100 +16200 = $69500 A moont allocated to howseacre = (6400 + 13